. In Exercise 16, we saw the following situation: Harada Construction has just bid on a major commercial project. The bid was based on a normal distribution representing Harada’s estimates of possible costs for the project. The distribution is centered on $6.5 million and has a standard deviation of $1.1 million. According to Harada’s estimates:

a. There is a 60% chance that costs will be less than $_______ million.

b. There is a 1% chance that the cost will be greater than $____ million.

c. There is a 45% chance that costs will be between $6.5 million and $______ million. (Make $6.5 million the lower boundary for your interval.)
Answer #1


Let X be the possible costs X follows normal distribution mean μ and standard deviation σ Given H $6.5 million ơ-$1.1 millionb) To find X such that P(X>X 0.01 1% chance P(X> X) 1 P(X< X) 1 0.01 P(X-X) = 0.99 Converting to z-scores P(Xc X) = P(Z-z) = 0.99 where 0.99 From standard normal tables or Excel function NORM.S.INV, we find the Z such that P(Z <z) 0.99 z NORM.S.INV(0.99) z 2.3263 Thus X-6.5- 2.3263 1.1 X 2.3263 1.16.5 X 9.0589 There is 1% chance that cost will be greater than $9.0589 million c) To find X such that P(6.5 < X < X) = 0.45 P(6.5 <X<X P(X <X) P(X < 6.5) 0.45 P(X < X) P(X< 6.5) Since $6.5 million is the mean, we know P(X < 6.5)- 0.5 0.45 P(X < X) - 0.5 P(X <X) 0.95 Converting to z-scores P(X <X)P(Z <z) 0.95 where 1% chance 1.1 From standard normal tables or Excel function NORM.S.INV, we find the Z such that P(Z <z) 0.95 z NORM.S.INV(0.95) z 1.645 Thus X-6.5- 1.645 1.1 X 1.645 1.1 6.5 X 8.3095 There is 45% chance that cost will be between $6.5 million and $8.3095 million